If the roots of quadratic equation px²-9x+4=0 are 4 and ½.Calculate the value of p.
Answer by Theo(13342)

roots are 4 and 1/2.
factors are (x-4) * (x-.5) = 0
multiply those factors together and you get x^2 - 4.5x + 2 = 0
multiply both sides of that equation by 2 and you get 2x^2 - 9x + 4 = 0
looks like p needs to be equal to 2.

your equation is 2x^2 - 9x + 4 = 0
factor that equation to get (2x-1) * (x-8) = 0
solve for x to get x = 1/2 and x = 8.

solution is good.